A 350kg motorcycle is moving at 20m/s. It then speeds up to 30m/s. How much work did the engine do to accelerate? *
5 points
1
Expert's answer
2020-11-23T10:28:19-0500

Let "t" be the time during which motorcycle was being accelerated. Then, the acceleration is:


"a = \dfrac{\Delta v}{t}"

where "\Delta v =v_f-v_i = 30m/s - 20m/s = 10m/s" is the velocity increase (difference between final "v_f" and initial "v_i" velocities). Then the engin force will be (according to the Newton's second law):


"F = ma = m\dfrac{\Delta v}{t}"

where "m = 350kg" is the mass of the motorcycle.

The distance travelled in time "t" is:


"d = v_0t + \dfrac{at^2}{2} = v_0t + \dfrac{\Delta vt}{2} = v_0t + \dfrac{\Delta vt}{2}"

The work done by the engin is:


"A = Fd\\
A = m\dfrac{\Delta v}{t}\left( v_0t + \dfrac{\Delta vt}{2} \right) = m\Delta v\left( v_0 + \dfrac{\Delta v}{2} \right)"

Substituting the numbers, obtain:

"A = 350\cdot 10\cdot \left( 20 + \dfrac{10}{2} \right) = 87500\space J"

Answer. 87500 J.